calculate the wavelenght in nm associated with proton moving at 1*103 m/s .

Grade:11

1 Answers

Shailendra Kumar Sharma
6 years ago
De Broglie equation
lamda =h/P
lambda = h/mv 
lambda = {6.63X10-34 }/{1.67X10-27 X 1 X 103}= .397X10-9 nm =.4nm so the wavelength is .4nm